A fair die is rolled till all the numbers 1, 2, ..., 6 appear on the top. What is the probability that at least 7 rolls are needed?


1
Expert's answer
2021-04-15T07:01:05-0400

"\frac{6\times6!+15\times6!}{6^7} = \frac{21\times6!}{6^7} = 0.054 =" 5.4%